# Calibrated and uncertain? Evaluating uncertainty estimates in binary classification models

### v1

@@ A.1 Neural networks with stochastic gradient descent
A simple approximation of the posterior distribution of variational inference may be obtained by assuming that only the value $\boldsymbol{\theta}=\boldsymbol{\theta}_{\text{max}}$ which maximizes the posterior $p(\boldsymbol{\theta}|\mathcal{D_{\text{train}}})$ contributes significantly to the integral in Equation 4. This assumption holds if there is just one mode, which is close to the mean and most of the probability mass is centered around it. This is the case if we assume the posterior to be approximately Gaussian. The maximum a posteriori estimate may be found by any optimization method which can maximize the posterior with respect to $\boldsymbol{\theta}$. In the case of deep learning we use stochastic gradient descent derived optimization techniques for expectation maximization with regularization, which may be interpreted in a Bayesian way as sampling from this assumed Gaussian-like posterior distribution as defined by the NN architecture and the stochastic gradient descent algorithm and hyperparameters (Duvenaud et al., 2016).
All deep learning algorithms were implemented in pytorch.
@@ A.2 Neural Network ensembles
In neural network ensemble models we independently train $N_{\text{E}}$ number of ensembles to get a set of samples $\boldsymbol{\hat{\theta}_{\text{max},i}}$ from the posterior distribution and use these to compute the estimated probability and uncertainty using Equations 4 and 5 as
@@ A.3 Conflictual loss
Here $i$ is the index of the network in the ensemble, $b$ is the index of the class that network is biased towards and $\beta$, the bias weight, is a hyperparameter which governs the strength of the bias.
The cross entropy loss of N data points is given by
$\text{CE}(\mathcal{D}_{N},\boldsymbol{\theta})=-\sum_{i=1}^{N}\sum_{j=1}^{k}c^{j}_{i}(\log Q(c^{j}=1|\boldsymbol{\theta},\mathbf{x_{i}}).$
@@ A.4 Monte Carlo Dropout
In deep learning using Monte Carlo Dropout (MCD) for uncertainty estimation (Gal and Ghahramani, 2016), the “dropping out” of weights commonly used during training to prevent overfitting, is applied at both training and inference phases. Specifically, during each forward pass through the network, random weights are set to zero according to a specified hyperparameter referred to as the dropout rate. MCD is related to ensemble learning. By performing multiple forward passes with different dropout masks, we generate a distribution of networks with associated predictions for a given input. Using a neural network with one infinitely wide hidden layer and MCD can theoretically be shown to be equivalent to a Gaussian Process, and therefore MCD is frequently referred to as approximate Bayesian inference. The estimated MCD probabilities and uncertainties presented in this study are the sample mean and standard deviations of 500 forward passes through the network.
@@ A.5 Evidential deep learning
where the components of $\nu$ belong to the standard k1k-1 simplex and $\text{B}(\boldsymbol{\alpha})$ is the multivariate beta function. $\alpha^{j}=e^{j}+1$ where $e^{j}$ is the “evidence”, in this case $e=\texttt{softplus}(x)=\log(1+\exp(x))$ of the logit outputs $\mathbf{f}(\mathbf{x},\boldsymbol{\theta})\in\mathbf{R}^{k}$ of the last non-softmax layer of the deep neural network.
@@
where $S_{i}=\sum_{k=1}^{k}\alpha_{i}^{k}$ and $\tilde{\boldsymbol{\alpha}}=\mathbf{c}_{i}+(1-\mathbf{c}_{i})\cdot\boldsymbol{\alpha}(\mathbf{x}_{i})$. This is the sum of the Bayes risk and the weighted KL-divergence between the estimated distribution and the uniform Dirichlet distribution. The weight $\lambda(t)$ determines the strength of the second term, and is usually implemented as a linear function of the current epoch during training. In our case $\lambda(t)=\lambda_{0}t/200$ and $\lambda_{0}$ is a hyperparameter. The training algorithm aims to ensure that data points with no evidence are assigned a low belief mass, resulting in high uncertainty.
@@
A.6 Nonparametric Bayesian Machine Learning
Nonparametric Bayesian inference algorithms provide principled methods of predictive inference for complex data. The connection between nonparametric Bayesian statistics and deep learning provides a bridge between the disciplines of statistics and machine learning and allows us to approach the latter with a well developed theoretical framework. We have included two Bayesian nonparametric inference algorithm in this study to show the similarities and differences to the approximate deep learning algorithms.
@@ A.7 Gaussian Process
A Gaussian Process (GP) is a Bayesian nonparametric approach to machine learning and closely related to deep learning (Rasmussen and Williams, 2006). In a GP model, we assume the relation $y=f(x)+\epsilon_{y}$ where $y$ is the variate we want to predict, $x$ are the rest of the variates, $f(x)$ is some function and $\epsilon_{y}$ is normally distributed noise. This makes $f(x)=y-\epsilon_{y}$ also normally distributed as $\mathcal{N}(\mu_{f},\sigma_{f})$, with $\mu_{f}$ and $\sigma_{f}$ given by the kernel function $\kappa(x,x_{*})$ whose value is dependent on the relation between the training data points $x$ and the test point(s) $x_{*}$. In a binary classification task, the GP outputs the mean and variance of the logits $f_{*}=f(\mathbf{x_{*}})$ for each test point $x_{*}$ which we may transform to have the properties of a binomial probability distribution through the sigmoid function to perform logistic regression. If we use a kernel function which diverges for $x$ far from $\mu_{x}$, such as the radial basis function kernel, the estimated class probabilities converge towards $1/2$ far from the data and the estimated uncertainties also converge towards some value decided by the implicit prior given by the chosen kernel and hyperparameters. To keep the mathematical framework consistent with the rest of the analysis and bypass the difficulty of evaluating the intractable integral caused by the sigmoid function, probabilities and standard uncertainties are approximated using $T=1000$ Monte Carlo samples generated from $\mathcal{N}(\mu_{f*},\sigma_{f*})$  [^3] The difference between these probability estimates for our test set and other approximations of the class probabilities such as the approximation by David J.C. MacKay (MacKay, 1992) or the implementation used in Scikit taken from (Williams and Barber, 1998) is less than 0.01.:
$\mathbb{E}_{f_{*}}[\texttt{sigmoid}(f_{*})]=\int\texttt{sigmoid}(f_{*})p(f_{*}|\mathcal{D}_{\text{train}})df_{*}$
$\approx\frac{1}{T}\sum_{t=1}^{T}\texttt{sigmoid}(f_{*,t})=\bar{Q}_{\text{GP}}(\mathbf{c}|\mathbf{x},\mathcal{D}_{\text{train}})$
$u_{GP}(\mathbf{c}|\mathbf{x},\mathcal{D}_{\text{train}})^{2}=\frac{1}{T}\sum_{t=1}^{T}(\texttt{sigmoid}(f_{*,t})-\bar{Q}_{\text{GP}}(\mathbf{c}|\mathbf{x},\mathcal{D}_{\text{train}}))^{2}$
with $p(f_{*}|\mathcal{D})=\mathcal{N}(\mu_{f*},\sigma_{f*})$.
We used a modified version of the GaussianProcessClassifier class in scikit learn (Pedregosa et al., 2011) to extract the mean and standard deviation of the latent function posterior distribution. A radial basis function kernel with initial output variance 2 and length scale 8 was used. These parameters are automatically optimized and are given in Table1(b).
@@ A.8 Dirichlet Process Mixture Model
The last model in the study is a Dirichlet process mixture model (DPMM) of product kernels. The Dirichlet Process is a stochastic process which generalizes the Dirichlet distribution, the conjugate prior for a discrete nominal variate with a fixed number of categories, into the prior for infinitely many categories, making the mixture model nonparametric (Li et al., 2019). Unlike the GP, this framework makes no simplifying Gaussian assumptions, and should be able to model more complex distributions. The algorithm is also the only one in the study that approximates the full joint distribution $p(\mathbf{c},\mathbf{x}|\mathcal{D})$. This makes the data model powerful but the calculations become more difficult and require extensive resource demanding Markov-Chain Monte Carlo simulations. The predictive posterior probability for a new data point is given by
@@
For our study we used the implementation in v0.2.1 of the R-package inferno (Porta-Mana et al., 2024) [^4] https://github.com/pglpm/inferno/ based on the mathematical framework of (Dunson and Bhattacharya, 2011; Ishwaran and Zarepour, 2002), which is a flexible framework for regression and classification on manifolds. The prior distribution and hyperparameter choices are kept default as described in the accompanying documentation (Porta-Mana, 2025). The most relevant hyperparameter choice is the use of a Bernoulli distribution with a $\text{Beta}(1,1)$ prior for the class variate. Changing the value of the Beta hyperparameters corresponds to changing the strength of our “noninformative” belief of equal probability in each class. 1200 samples of $\mathbf{w},\boldsymbol{\alpha}$ for the approximation of the joint posterior distribution were collected using Gibbs sampling and 10 Markov-Chain Monte Carlo chains with a burn-in of 3600 samples and early stopping (Gong and Flegal, 2016). We approximate the expected value and standard uncertainty of the conditional class probability by the sample mean and standard deviation of the conditional class probabilities calculated from the Monte Carlo samples.
$\bar{Q}_{\text{DPMM}}(\mathbf{c}|\mathbf{x},\mathcal{D}_{\text{train}})=\frac{1}{T}\sum_{t=1}^{T}Q(\mathbf{c}|\mathbf{x},\mathbf{w}_{t},\boldsymbol{\alpha}_{t})$
$u_{\text{DPMM}}^{2}(\mathbf{c}|\mathbf{x},\mathcal{D}_{\text{train}})=\frac{1}{T}\sum_{t=1}^{T}(Q(\mathbf{c}|\mathbf{x},\mathbf{w}_{t},\boldsymbol{\alpha}_{t})-\bar{Q}_{\text{DPMM}}(\mathbf{c}|\mathbf{x},\mathcal{D}_{\text{train}}))^{2}$
@@ Appendix B Neural network hyperparameters
Figure 7: Results for different hyperparameter settings for neural network ensemble (NNE). The columns contain the results for learning rates 0.0001, 0.001 and 0.01. The color of the bars indicates the depth and width of the hidden layers of the network. Small, medium and large has 1, 3, and 8 hidden layers with 20, 200 and 2000 nodes respectively. The top row plots shows the accuracy on the validation set, with the green dashed line indicates the optimal accuracy for that validation set. The second row plots shows the Wasserstein distance (WD) of the validation set. The two last rows of plots show the average estimated class probability and uncertainty for out-of-distribution test points. The error bars indicate the 2.5-97.5 percentiles.
@@
The neural networks we have used in the study are all fully connected neural networks with ReLU activation functions. Changing the activation function does seem to change the extrapolation values of the estimated probabilities and uncertainties, but we choose to focus on the ReLU activation function as it is the most widely used in deep learning. They are trained using stochastic gradient descent with the AdamW modification of the Adam optimizer(Kingma and Ba, 2015) so we could control the weight decay. The models are trained on batches of $[128,128,256,256,1024,1024,2048]$ samples for $N_{\text{train}}=[250,500,1000,2000,3000,5000,10000]$. The models were each trained for a maximum of 200 epochs (substantially more than needed) with early stopping if the loss did not decrease within 20 epochs.
@@
A grid search was run over hyperparameter settings where we investigated the influence of changing learning rate (0.01, 0.001 and 0.0001), weight decay (0.1, 0.01 and 0.001) and network size (1, 3, 8 hidden layer(s) with 20, 200, 2000 nodes respectively) as well as the bias weight for CL (0.1, 0.2, 0.5), the annealing coefficient for EDL (0.0002/200, 0.0002/200, 0.02/200) and the drop out rate for MCD (0.1, 0.3, 0.5). For the ensembles, one ensemble with 20 networks was trained for each of the 81 different combinations of hyperparameters, while for the single network EDL and MCD algorithms 20 separate networks were trained for each combination. Weight decay was found to not have any major influence on our results, so for the study, weight decay was set to 0.01 for all algorithms. The accuracy and Wasserstein distance (WD) of the validation set, the average OOD estimated class probability and the average OOD uncertainty was calculated and are presented in Figures 7-10 as bar plots with error bars indicating the 2.5-97.5 percentiles. The $x$-axis of the plots show the results for different number of training points, while the color of the bar indicates the network size. The accuracy is a good measure of whether the network is properly trained. For dataset A, the optimal accuracy is 73.82% and for dataset B the optimal accuracy is 73.88%. Based on the results, any networks with less than 70% accuracy even with just 250 data points should be considered very poorly fitted. The WD is a measure of calibration, and for the validation set it should not be strongly influenced by any bias in the tails of the distribution. It is expected to go down with increasing number of training data, and hyperparameter settings which do not give results that follow this may be poor choices. The OOD estimated probability and uncertainty is also shown to see if the choice of hyperparameter would influence the results of the study. Some hyperparameter combinations do produce high uncertainties and more conservative estimated class probabilities, seemingly in conflict with our study results, but these tend to have lower accuracy scores and higher WD scores. The rest of the OOD results confirm our conclusion that NN-based estimated probabilities and uncertainties approach extreme probabilities and zero uncertainty.
@@
Appendix C Identically distributed classes
(a) Neural network ensemble
(b) Evidential deep learning
Figure 11: Results for different hyperparameter settings for neural network ensemble and evidential deep learning. The columns contain the results for learning rates 0.0001, 0.001 and 0.01. The color of the bars indicates the size and depth of the hidden layers of the network. Small, medium and large has 1, 3, and 8 hidden layers with 20, 200 and 2000 nodes respectively. The first and second rows of plots shows the mean and standard deviation of the estimated class probability for out-of-distribution test points. The bottom two rows of plots shows the mean and standard deviation of the estimated uncertainty for out-of-distribution test points. The error bars indicate the 2.5-97.5 percentiles.
As an additional test we created a dataset with two statistically identical classes by using the parameters of class 2 in dataset A when generating data for both classes. We ran the same hyperparameter grid search for all deep learning models. The results for NNE and EDL are presented in Figure 11. NNE, CL and MCD all produced class probability estimates which varied greatly both across hyperparameter values and as a function of polar angle. Evidential deep learning stood out as the only algorithm, which (for small enough $\lambda_{0}$) produced estimated probabilities of 0.5. Interestingly, the uncertainty was still close to zero. The GP and DPMM algorithms struggled with this dataset, especially for large numbers of data points so we could only evaluate performance on 2000 or less training data points. The GP hyperparameter optimization produced an output variance, which quickly approached zero as more data was added, leading to close to zero uncertainty everywhere, while the estimated probability stayed close to 0.5. For in distribution data, the DPMM gave probability estimates which were pretty noisy, as expected, but fluctuate around 0.5. Estimates for OOD data were approximately 0.5 and uncertainties were close to $0.5/\sqrt{3}$, in agreement with the chosen prior.
Appendix D 2D distributions

### v2

@@ A.1 Neural networks with stochastic gradient descent
A simple approximation of the integral in Equation 7 may be obtained by assuming that the only significant contribution to the integral comes from values of $\boldsymbol{\theta}$ close to the value $\boldsymbol{\theta}_{\text{max}}$ which maximizes the posterior $p(\boldsymbol{\theta}|\mathcal{D_{\text{train}}})$. This approximation is appropriate if there is just one mode, which is close to the mean and most of the probability mass is centered around it. This is the case if we assume the posterior to be approximately Gaussian. The maximum a posteriori estimate may be found by any optimization method which can maximize the posterior with respect to $\boldsymbol{\theta}$. In the case of deep learning we use stochastic gradient descent derived optimization techniques for expectation maximization with regularization, which may be interpreted in a Bayesian way as sampling from this assumed Gaussian-like posterior distribution as defined by the NN architecture and the stochastic gradient descent algorithm and hyperparameters (Duvenaud et al., 2016).
All deep learning algorithms in the study were implemented in pytorch. All deep learning classifiers are simple fully connected neural networks with a softmax-function as the last normalizing layer unless otherwise specified. Details on the chosen architectures and hyperparameter settings are presented in Appendix B.
@@ A.2 Neural Network ensembles
In neural network ensemble models we independently train $N_{\text{E}}$ neural networks to get a set of samples $\boldsymbol{\hat{\theta}_{\text{max},i}}$ from the approximate posterior distribution and use these to compute the estimated probability and uncertainty using Equations 7 and 8 as
@@
We set the number of ensembles to $N_{\text{E}}=20$ as in realistic cases it is usually not feasible to train many more ensembles than that. We tested varying $N_{\text{E}}$ between 5 and 50, and found that this made no significant difference to the results.
@@ A.3 Conflictual loss
Here $i$ is the index of the network in the ensemble, $b$ is the index of the class that network is biased towards and $\beta$, the bias weight, is a hyperparameter which governs the strength of the bias. The cross entropy loss of N data points is given by
$\text{CE}(\mathcal{D}_{N},\boldsymbol{\theta})=-\sum_{i=1}^{N}\sum_{j=1}^{k}c^{j}_{i}\log Q(c^{j}=1|\boldsymbol{\theta},\mathbf{x_{i}}).$
@@ A.4 Monte Carlo Dropout
In deep learning using Monte Carlo Dropout (MCD) for uncertainty estimation (Gal and Ghahramani, 2016), the “dropping out” of weights commonly used during training to prevent overfitting, is applied at both training and inference phases. Specifically, during each forward pass through the network, random weights are set to zero according to a specified hyperparameter referred to as the dropout rate. MCD is related to ensemble learning. By performing multiple forward passes with different dropout masks, we generate a distribution of networks with associated predictions for a given input. Using a neural network with one infinitely wide hidden layer and MCD can theoretically be shown to be equivalent to a Gaussian Process, and therefore MCD is frequently referred to as approximate Bayesian inference. The estimated MCD probabilities and uncertainties presented in this study are the sample mean and standard deviations of 200 forward passes through the network. Changing the number of forward passes to 50 or 500 instead had no significant effect on our results.
@@ A.5 Evidential deep learning
where the components of $\nu$ belong to the standard k1k-1 simplex and $\text{B}(\boldsymbol{\alpha})$ is the multivariate beta function. $\alpha^{j}=e^{j}+1$ where $e^{j}$ is the “evidence”, in this case $e^{j}=\texttt{softplus}(f^{j})=\log(1+\exp(f^{j}))$ of the logit outputs $\mathbf{f}(\mathbf{x},\boldsymbol{\theta})\in\mathbb{R}^{k}$ of the last non-softmax layer of the deep neural network.
@@
where $S_{i}=\sum_{k=1}^{k}\alpha_{i}^{k}$ and $\tilde{\boldsymbol{\alpha}}=\mathbf{c}_{i}+(1-\mathbf{c}_{i})\cdot\boldsymbol{\alpha}(\mathbf{x}_{i})$. This is the sum of the Bayes risk and the weighted KL-divergence between the estimated distribution and the uniform Dirichlet distribution. The annealing weight $\lambda(t)$ determines the strength of the second term, and is usually implemented as a linear function of the current epoch during training. In our case $\lambda(t)=\lambda_{0}t/200$ and $\lambda_{0}$ is a hyperparameter. The training algorithm aims to ensure that data points with no evidence are assigned a low belief mass, resulting in high uncertainty.
@@
A.6 Nonparametric Bayesian inference
Nonparametric Bayesian inference algorithms provide principled methods of predictive inference for complex data. The connection between nonparametric Bayesian statistics and deep learning provides a bridge between the disciplines of statistics and machine learning and allows us to approach the latter with a well developed theoretical framework. We have included two nonparametric Bayesian inference algorithm in this study to show the similarities and differences to the approximate deep learning algorithms.
@@ A.7 Gaussian Process
A Gaussian Process (GP) is a Bayesian nonparametric approach to machine learning and closely related to deep learning (Rasmussen and Williams, 2006). In a GP model, we assume the relation $y=f(x)+\epsilon_{y}$ where $y$ is the variate we want to predict, $x$ are the rest of the variates, $f(x)$ is some function and $\epsilon_{y}$ is normally distributed noise. This makes $f(x)=y-\epsilon_{y}$ also normally distributed as $\mathcal{N}(\mu_{f},\sigma_{f})$, with $\mu_{f}$ and $\sigma_{f}$ given by the kernel function $\kappa(x,x_{*})$ whose value is dependent on the relation between the training data points $x$ and the test point(s) $x_{*}$. In a binary classification task, the GP outputs the mean and variance of the logits $f_{*}=f(\mathbf{x_{*}})$ for each test point $x_{*}$ which we may transform to have the properties of a binomial probability distribution through the sigmoid function to perform logistic regression.
If we use a kernel function which diverges for $x$ far from $\mu_{x}$, such as the radial basis function kernel, the estimated class probabilities converge towards $1/2$ far from the data and the estimated uncertainties also converge towards some value decided by the implicit prior given by the chosen kernel and hyperparameters. To keep the mathematical framework consistent with the rest of the analysis and bypass the difficulty of evaluating the intractable integral caused by the sigmoid function, probabilities and standard uncertainties are approximated using $T=1000$ Monte Carlo samples generated from $\mathcal{N}(\mu_{f*},\sigma_{f*})$  [^5] The difference between these probability estimates for our test set and other approximations of the class probabilities such as the approximation by David J.C. MacKay (MacKay, 1992) or the implementation used in Scikit taken from (Williams and Barber, 1998) is less than 0.01.:
$\mathbb{E}_{f_{*}}[\texttt{sigmoid}(f_{*})]=\int\texttt{sigmoid}(f_{*})p(f_{*}|\mathcal{D}_{N})df_{*}$
$\approx\frac{1}{T}\sum_{t=1}^{T}\texttt{sigmoid}(f_{*,t})=\bar{Q}_{\text{GP}}(\mathbf{c}|\mathbf{x},\mathcal{D}_{N})$
$u_{GP}(\mathbf{c}|\mathbf{x},\mathcal{D}_{N})^{2}=\frac{1}{T}\sum_{t=1}^{T}(\texttt{sigmoid}(f_{*,t})-\bar{Q}_{\text{GP}}(\mathbf{c}|\mathbf{x},\mathcal{D}_{N}))^{2}$
with $p(f_{*}|\mathcal{D}_{N})=\mathcal{N}(\mu_{f*},\sigma_{f*})$.
We used a modified version of the GaussianProcessClassifier class in scikit learn (Pedregosa et al., 2011) to extract the mean and standard deviation of the latent function posterior distribution. A radial basis function kernel with initial output variance 2 and length scale 8 was used. These parameters are automatically optimized and are presented in Table 1(b). The use of a radial basis kernel with optimized parameters expresses a prior belief that we are looking for smooth functions of limited amplitude.
@@ A.8 Dirichlet Process Mixture Model
The last model in the study is a Dirichlet process mixture model (DPMM) of product kernels. The Dirichlet Process is a stochastic process which generalizes the Dirichlet distribution, the conjugate prior for a discrete nominal variate with a fixed number of categories, into the prior for infinitely many categories, making the mixture model nonparametric (Li et al., 2019). Unlike the GP, this framework relaxes the global Gaussian process assumptions and can model multi-modal or non-smooth structures. The algorithm is also the only one in the study that approximates the full joint distribution $p(\mathbf{c},\mathbf{x}|\mathcal{D})$. This makes the data model powerful but the calculations become more difficult and require extensive resource demanding Markov-Chain Monte Carlo simulations. The predictive posterior probability for a new data point is given by
@@
For our study we used the implementation in v0.2.1 of the R-package inferno (Porta-Mana et al., 2024) [^6] https://github.com/pglpm/inferno/ based on the mathematical framework of (Dunson and Bhattacharya, 2011; Ishwaran and Zarepour, 2002), which is a flexible framework for regression and classification on manifolds. The prior distribution and hyperparameter choices are kept default as described in the accompanying documentation (Porta-Mana, 2025). For the class variate, we used a mixture of Bernoulli distributions with a $\text{Beta}(1,1)$ prior, and for the input features we used a mixture of axis-aligned Gaussian distributions with prior hyperparameters automatically set by the data. Changing the value of the Beta hyperparameter $c$ corresponds to changing the strength of our “noninformative” belief of equal probability in each class. 1200 samples of $\mathbf{w},\boldsymbol{\alpha}$ for the approximation of the joint posterior distribution were collected using Gibbs sampling and 10 Markov-Chain Monte Carlo (MCMC) chains with a burn-in of 3600 samples and early stopping (Gong and Flegal, 2016). Issues related to MCMC convergence and truncation can cause numerical artifacts in these kind of models, so it is important to inspect both the MCMC diagnostics results and the inference results carefully. We approximate the expected value and standard uncertainty of the conditional class probability by the sample mean and standard deviation of the conditional class probabilities calculated from the Monte Carlo samples.
$\bar{Q}_{\text{DPMM}}(\mathbf{c}|\mathbf{x},\mathcal{D}_{N})=\frac{1}{T}\sum_{t=1}^{T}Q(\mathbf{c}|\mathbf{x},\mathbf{w}_{t},\boldsymbol{\alpha}_{t})$
$u_{\text{DPMM}}^{2}(\mathbf{c}|\mathbf{x},\mathcal{D}_{N})=\frac{1}{T}\sum_{t=1}^{T}(Q(\mathbf{c}|\mathbf{x},\mathbf{w}_{t},\boldsymbol{\alpha}_{t})-\bar{Q}_{\text{DPMM}}(\mathbf{c}|\mathbf{x},\mathcal{D}_{N}))^{2}$
@@ Appendix B Neural network hyperparameters
Figure 9: Results for different hyperparameter settings for neural network ensemble (NNE). The columns contain the results for learning rates 0.0001, 0.001 and 0.01. The color of the bars indicates the depth and width of the hidden layers of the network. Small, medium and large has 1, 3, and 8 hidden layers with 20, 200 and 2000 nodes respectively. The top row plots shows the accuracy on the validation set, with the green dashed line indicates the LRFD accuracy for that validation set. The second row plots shows the Wasserstein distance (WD) of the validation set. The two last rows of plots show the average estimated class probability and uncertainty for out-of-distribution test points. The error bars indicate the 2.5-97.5 percentiles.
@@
The neural networks we have used in the main study are all fully connected neural networks with ReLU activation functions. They are trained using stochastic gradient descent with the AdamW modification of the Adam optimizer (Kingma and Ba, 2015). The models are trained on batches of $[128,128,256,256,1024,1024,2048]$ samples for $N_{\text{train}}=[250,500,1000,2000,3000,5000,10000]$. The models were each trained for a maximum of 200 epochs (substantially more than needed) with early stopping if the loss did not decrease within 20 epochs. The early stopping criteria, as well as the way the hyperparameters are chosen as described below, biases the results towards networks with good global calibration properties and does not take OOD performance into account.
A grid search was run over hyperparameter settings where we investigated the influence of changing learning rate (0.01, 0.001 and 0.0001), weight decay (0.1, 0.01 and 0.001) and network size (1, 3, 8 hidden layer(s) with 20, 200, 2000 nodes respectively) as well as the bias weight for CL (0.1, 0.2, 0.5), the annealing coefficient for EDL (0.0002/200, 0.0002/200, 0.02/200) and the drop out rate for MCD (0.1, 0.3, 0.5). For the ensembles, one ensemble with 20 networks was trained for each of the 81 different combinations of hyperparameters, while for the single network EDL and MCD algorithms 20 separate networks were trained for each combination. Weight decay was found to not have any major influence on our results, so for the study, weight decay was set to 0.01 for all algorithms.
@@
The accuracy and Wasserstein distance (WD) of the validation set, the average OOD-far estimated class probability and the average OOD-far uncertainty was calculated and are presented in Figures 9-12 as bar plots with error bars indicating the 2.5-97.5 percentiles over different runs, including the runs with varying weight decay. The $x$-axis of the plots show the results for different number of training points, while the color of the bar indicates the network size. The accuracy is a good measure of whether the network is properly trained. For dataset A, the optimal accuracy of the validation set, if the long run frequency was known, is 73.82% and for dataset B the accuracy is 73.88%. Based on the results, any networks with less than 70% accuracy even with just 250 data points should be considered very poorly fitted.
The WD is a global measure of calibration, and for the validation set it should not be strongly influenced by any bias in the tails of the distribution. It is expected to go down with increasing number of training data, and hyperparameter settings which do not give results that follow this trend are considered to be poor choices. The OOD estimated probability and uncertainty is also shown to see if the choice of hyperparameter would influence the results of the study. Some hyperparameter combinations do produce high uncertainties and more conservative estimated class probabilities, seemingly in conflict with our study results, but these tend to have lower accuracy scores and higher WD scores. The rest of the OOD results confirm our conclusion that the NN-based estimated probabilities and uncertainties for our dataset approach extreme probabilities and zero uncertainty.
@@
We also checked the effect of different hyperparameter choices for the DPMM algorithm. We tested that the estimated OOD-far uncertainties were approximately equal to the predictions of equation 5 for $c=[0.1,0.5,1,2]$. Lower values for c give a slower convergence (as a function of $r$) towards the theory predictions and vice versa. All values for $c$ showed an increase in the predicted class 1 probability estimates towards 1 for larger values of $N_{\text{train}}$. Class 1 is more dominant in the tails of the in-distribution data and may therefore be assigned higher probabilities that should in principle be balanced by the rest of the components in the mixture, but in practice are not. Inspections of the trace plots and diagnostics do not indicate that these results are due to MCMC convergence issues.
@@
Appendix C Changing the activation function
(a) Dataset A.
(b) Dataset B.
Figure 13: Estimated probabilities (top row) and uncertainties (bottom row) for class 1 for the different algorithms for the two datsets as a function of radius $|\mathbf{x}|$. The error bars indicate the entire spread of the data over polar angle $\phi$, while the markers indicate the sample average. The dashed black line in the plots in the bottom rows indicates the standard deviation of the predictive posterior of a binomial model with $N=0$ and prior $\text{Beta}(1,1)$.
Figure 14: Top: Average estimated out-of-distribution probabilities (top rows) and uncertainties (bottom rows) as a function of $N_{\text{train}}$ for the test set (left), the OOD-near test set (middle) and the OOD-far test set (right) with the deep learning models using the tanh-activation function. The error bars indicate the entire spread of the data over polar angle $\phi$. The colors are similar to Figure 5. The dashed black line in the plots in the bottom rows indicates the standard deviation of the predictive posterior of a binomial model with $N=0$ and prior $\text{Beta}(1,1)$.
We also tried changing the activation function in the neural networks to avoid the known pathologies of ReLU neural networks. Using the tanh activation function gave results that were qualitatively different from the previous results. Figure 13 shows the average and spread in class probability and uncertainty estimates over $\phi$ as a function of $r$ for values of $r$ ranging from 0 to 1000. In this case, the estimated probabilities stabilize around 0.8 instead of 1 and uncertainties are not zero. For dataset A there is no distinct increase in uncertainty for OOD data for NNE, CL and MCD, but for EDL there is an observed increase. However, the estimates vary wildly across $\phi$ and would not be reliable in a practical setting. For dataset B we see a profound increase in uncertainties as a function of increasing $r$ for NNE and CL for 5000 training data points. The trend holds for $N_{\text{train}}\geq 1000$ as can be seen in Figure 14. The estimates still do not give class probabilities of 0.5, and uncertainties are much lower than DPMM and GP for the OOD-near test points. EDL estimates behave very unpredictably for this dataset while MCD has similar behavior as for dataset A.
Appendix D Reliability diagrams
(a) Dataset A
(b) Dataset B
Figure 15: Reliability diagrams calculated using the test datasets over 15 uniform bins in $[0,1]$ for the two datasets. The the x-axis indicates the estimated class probability of class 1, while the y-axis of the diagram indicates the empirical frequency of class 1 of the bin. The top plots for each dataset are the results for $N_{\text{train}}=500$ and the bottom for $N_{\text{train}}=5000$. Under each reliability plot is a histogram showing the number of data points in the test set with estimated class probabilities belonging to that bin.
Reliability diagrams are a common tool for evaluating calibration of class probabilities. The reliability diagrams for a selection of the trained models are presented in Figure 15. They were calculated using 15 uniform bins over $[0,1]$. The plots quantitatively support the claim that all models are well calibrated outside the tail of the distributions. For the bins close to 1, it can be seen that the deep learning models are overconfident with respect to the calculated empirical frequency distribution for dataset A, but not for dataset B.
Appendix E 2D distributions